a sphere of radius r moving with a velocity v under streamline conditions in a viscous fluid experiences a retarding force, F, given by F = Krv where K is constant. Derive the S.I units of K in terms of the base units.
1
Expert's answer
2019-11-04T09:54:18-0500
"[F]=[K][r][v]"

"kg\cdot m\cdot s^{-2} =[K](m)(m\cdot s^{-1})"

"[K]=kg\cdot m^{-1}\cdot s^{-1}"